Calculate ΔH for this reaction: CH4(g) + NH3(g) ---> HCN(g) + 3H2(g)
given:
N2(g) + 3H2(g) ---> 2NH3(g)
ΔH = −91.8 kJ
C(s) + 2H2(g) ---> CH4(g)
ΔH = −74.9 kJ
H2(g) + 2C(s) + N2(g) ---> 2HCN(g)
ΔH = +270.3 kJ
1
Expert's answer
2020-11-23T06:58:49-0500

CH4 : b reversible

CH4(g) "\to" C(graphite) + 2H2(g)

H = + 74.9 KJ


NH3 : a÷2 +reversible

NH3(g) "\to" "\frac{1}{2}" N2(g) + "\frac{3}{2}" H2(g)

H = + 45.9 KJ


HCN : c÷2 non reversible

"\frac{1}{2}" H2(g) +C(graphite) + "\frac{1}{2}" N2(g) "\to" HCN(g)

H = +135.15KJ


CH4(g) + NH3(g) "\to" HCN(g) +3H2(g)

H = +74.9 KJ + 45.9KJ + 135.15KJ

H = +255.96KJ or

H = +256KJ